A customer calls the customer service of the e-commerce site where he is shopping. Customer service cannot be reached as it is full 60% of the time. How likely is a customer who makes 7 calls in a row to be connected to customer service on any one call?


1
Expert's answer
2021-06-18T07:54:22-0400

Binomial distribution: "P(X) = nCx p^x q^{n-x}"

Number of trials, n = 7

P(issue resolved in one hour), p = 0.60

"q = 1 - p = 0.40"

P( 7 calls in a row) = "P(7)"

"= 7C1*0.6^7*0.4^1"

= 7*0.01119744

=0.07838208